OpenStudy (meanhacker):

help solve the eigenvalue problem: y''+λ²y=0, y'(0)=0, y(π)=0

OpenStudy (anonymous):

so\[y"=-\lambda ^{2}\] integrating both sides we get\[\int\limits_{?}^{?}y"=-\int\limits_{?}^{?}\lambda ^{2}dx\] so we get \[y'=-\lambda ^{2}x+c\] since y'(0)=0...c=0.. so\[y'=-\lambda ^{2}x\] similarly proceed..integrate this 1st order equation..put in the given value of y and x...find c... and you will get the required general solution..

OpenStudy (blacksteel):

RaGhavv, I suspect your answer is incorrect, since I assume you meant \[y'' = -\lambda^2y\]

OpenStudy (anonymous):

\[y''+\lambda^2y=0\] gives the characteristic equation \[r^2+\lambda^2=0\] with roots \[r=\pm\lambda i\] So the solution is of the form \[y=C_1\sin\lambda x+C_2\cos \lambda x\] Use the given boundary conditions to solve for \(C_1,C_2\).

OpenStudy (meanhacker):

thank you @SithsAndGiggles

OpenStudy (anonymous):

You're welcome!

OpenStudy (meanhacker):

check out this grapher for windows by the way http://www.walterzorn.de/en/grapher/grapher_app.htm

OpenStudy (anonymous):

@blacksteel ya i wrongly thought the ques to be...\[y"+\lambda ^{2}=0\]